OpenStudy (anonymous):

Rewrite sin(2sin^-1x)

OpenStudy (jdoe0001):

sin(2sin^-1x) sin(2sin^-1x) <--- REwritten =)

OpenStudy (anonymous):

could you explain how you got that?

OpenStudy (jdoe0001):

easy, I just wrote it twice =), you asked to REWRITE, so I wrote it once, and rewrote it again

OpenStudy (anonymous):

Omg, lol.

OpenStudy (jdoe0001):

what are you supposed to rewrite it as specifically?

OpenStudy (anonymous):

Rewrite it in an algebraic expression in x

OpenStudy (raden):

use the double formula : sin2a = 2sinacosa

OpenStudy (jdoe0001):

hmmmm shoot.. I didn't read it well... I see the issue now =)

OpenStudy (anonymous):

I know but I don't know how to plug it in. It comes out to be like sqrt of 1-x^2 and i don't get how it comes out to be that way.

OpenStudy (raden):

so, sin(2sin^-1x) = 2sin(sin^-1 x) cos(sin^-1 x) next, we knowed that the identity of : cos^2 x = 1 - sin^2 x, then cosx = sqrt( 1 - sin^2 x) now we have : 2sin(sin^-1 x) cos(sin^-1 x) = 2sin(sin^-1 x) sqrt(1 - (sin(sin^-1 x)^2)) = 2 x sqrt(1-x^2)

OpenStudy (anonymous):

Thank you so much.

OpenStudy (raden):

you're welcome